OpenStudy (anonymous):

define y as a function of f 8x^2+7y^2-1

OpenStudy (anonymous):

You're typically given a function f(x) = 2x*3+5 (example) right? now thats the same that saying y=2x*3+5 now your asked to form a function f(y), so...

OpenStudy (anonymous):

so, I do y=x^2+7y^2-1

OpenStudy (anonymous):

that would be f(x,y)=x^2+7y^2-1 u have to go 8x^2+7y^2-1=0 8x^2=1- 7y^2 x^2=(1-7y^2)/8 x=sqrt((1-7y^2)/8) f(y)=sqrt((1-7y^2)/8) I think that's what ur asked to do :D

OpenStudy (anonymous):

thank you...

OpenStudy (anonymous):

ur welcome